Cranston Cranberry Cooperative (CCC) processes cranberries that are harvested in the local area. Barrels of cranberries arrive on trucks to CCC’s facility at a rate of 150 barrels per hour and are processed continuously at a rate of 100 barrels per hour. Trucks arrive at a uniform rate over 8 hours, from 6:00 a.m. until 2:00 p.m. What is the maximum number of barrels of cranberries that are waiting on the trucks at any given time? A. 100 barrels B. 200 barrels C. 400 barrels D. Infinitely many barrels
Answer #1

Correct Answer:

C

The additional barrels in waiting per hour = 150-50 = 50

So, number of barrels waiting at trucks in 8 hrs. = 8*50 = 400 numbers of barrels